OpenStudy (blyue):

Can anyone lend a hand at finding the range of a function? :) So.. I have this problem. f(x) = 1/x-7 + 2 I've tried to get this on my own, but I just can't. I feel silly. Especially since it seems like such an easy concept! I know how to find the domain alright, but the range just puzzles me. This is a question from a practice test, and I have the answers, but I can't find my way to get it... The answer is that: Domain: All real numbers except 7 Range: All real numbers except 2 So, would anyone mind lending a hand? Any help is appreciated. I'd just really like to pin down how to find the range once and for all. :)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f(x)=\frac{1}{x-7}+2 }\)

OpenStudy (solomonzelman):

Domain of the function := all x-values that you are capable of plugging into the function Range of the function := all y-values that can possible get (as an output) from the function

OpenStudy (solomonzelman):

Range will depend on the domain.

OpenStudy (solomonzelman):

You probably know that \(\color{#000000 }{ \displaystyle \frac{ 1 }{x-7} \ne0 }\) Because there isn't an x-value that would satisfy this equation. And therefore, \(\color{#000000 }{ \displaystyle \frac{ 1 }{x-7}+2 \ne2 }\) So, you can see that f(x) is never equal to what number? (Note, that you can possible get absolute any output, besides that one number that you may never get,)

OpenStudy (blyue):

You sir, are magic. I think I finally get it. Thank you so much for the help, seriously! :D

OpenStudy (solomonzelman):

You are welcome!